Andy is four times as old as Wendy . In twelve years Wendy's age will be half of Andy's.find their ages now.
Answer by algebrahouse.com(1659)

"Andy is four times as old as Wendy . In twelve years Wendy's age will be half of Andy's.find their ages now."

x = Wendy's age now
4x = Andy's age now

x + 12 = Wendy's age in 12 years
4x + 12 = Andy's age in 12 years

x + 12 = (1/2)(4x + 12) {in twelve years, Wendy's age will equal half Andy's age}
x + 12 = 2x + 6 {used distributive property}
12 = x + 6 {subtracted x from both sides}
x = 6 {subtracted 6 from both sides}
4x = 24 {substituted 6, in for x, into 4x}

Wendy is 6 now and Andy is 24 now
www.algebrahouse.com